OpenStudy (anonymous):

Need Algebra Help Please. Will Fan & Medal :) Factor each expression. 44. 6k^2 - 10kl + 4l^2 = 54. -2h^2 + 4h + 70 = 68. w^2 + 24w + 144 = 70. 25x^2 - 36 =

OpenStudy (anonymous):

@jim_thompson5910 Can you help me please?

OpenStudy (anonymous):

@KendrickLamar2014

jimthompson5910 (jim_thompson5910):

`44. 6k^2 - 10kl + 4l^2` the first coefficient is 6 the last coefficient is 4 multiply them: 6*4 = 24 now ask yourself: what two numbers multiply to 24 and add to -10 (middle coefficient)?

OpenStudy (anonymous):

@jim_thompson5910 Negative 6 & positive 4?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

it should be -6 and -4

OpenStudy (anonymous):

44. 6k^2 - 10kl + 4l^2 the first coefficient is 6 the last coefficient is 4 multiply them: 6*4 = 24 now ask yourself: what two numbers multiply to 24 & add to -10 (middle coefficient)? So finding the coefficients are steps to factor the expression?

jimthompson5910 (jim_thompson5910):

Use that fact to break up the `-10kl` into `-6kl-4kl` then factor by grouping \[\Large 6k^2 - 10kl + 4l^2\] \[\Large 6k^2 - 6kl - 4kl + 4l^2\] \[\Large (6k^2 - 6kl) + (-4kl + 4l^2)\] \[\Large 6k(k - l) + (-4kl + 4l^2)\] \[\Large 6k(k - l) -4l(k - l)\] \[\Large (6k - 4l)(k - l)\] \[\Large (2(3k - 2l))(k - l)\] \[\Large 2(3k - 2l)(k - l)\]

OpenStudy (anonymous):

So 2(3k - 2l) (k - l) is the first answer?

jimthompson5910 (jim_thompson5910):

yeah

OpenStudy (anonymous):

Can you help me with the rest?

OpenStudy (anonymous):

54. -2h^2 + 4h + 70 =

jimthompson5910 (jim_thompson5910):

the site is being very slow

jimthompson5910 (jim_thompson5910):

you'll follow the same steps for the rest -2*70 = -140 find two numbers that multiply to -140 and add to 4

OpenStudy (anonymous):

Can you walk me threw it please?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

`find two numbers that multiply to -140 and add to 4` tell me what those two numbers are

OpenStudy (anonymous):

-2 times 70 = -140

OpenStudy (anonymous):

plus 4 = -144

OpenStudy (anonymous):

@jim_thompson5910 So -144?

jimthompson5910 (jim_thompson5910):

what two whole numbers multiply to -140 ?

OpenStudy (anonymous):

2 and 70

jimthompson5910 (jim_thompson5910):

2*70 = 140 2 + 70 = 72 we need the two factors to add to 4 (not 72)

OpenStudy (anonymous):

I really dont get it.

jimthompson5910 (jim_thompson5910):

this video might be able to help https://www.youtube.com/watch?v=mNrk3TTYnS8 it offers some good examples